OpenStudy (anonymous):

5/(6-4sqrt7)-(8-5sqrt7)

OpenStudy (anonymous):

\[\frac{ 5 }{ (6-4√7)-(8-5√7) }\]

OpenStudy (anonymous):

Simplify or solve? If solve..just use a calculator

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

I got: \[\frac{ 10+5√7 }{ 3 }\]

OpenStudy (aravindg):

can you show how you got that ?

OpenStudy (anonymous):

distribute the minus sign in the denominator and simp the bottom first.

mathslover (mathslover):

By the way , you are correct @lornbeach . EXCELLENT work!

OpenStudy (mertsj):

You are correct, lornbeach

OpenStudy (anonymous):

thank youu (:

OpenStudy (mertsj):

yw

mathslover (mathslover):

You're welcome.

mathslover (mathslover):

By the way check if you did the same : \[\cfrac{5}{6 - 4\sqrt{7} \color{blue}{-8 + 5\sqrt{7}}}\] \(\cfrac{5}{\color{blue}{6-8} \color{red}{ + 5\sqrt{7} - 4\sqrt{7}}}\) \(\cfrac{5}{-2 + \sqrt{7}}\) \(\cfrac{5 \color{orange}{(\sqrt{7} +2)}}{(-2 + \sqrt{7})\color{orange}{(\sqrt{7} + 2)}}\) \(\cfrac{10 + 5\sqrt{7}}{7-4}\) \(\cfrac{10 + 5\sqrt{7}}{3}\)